OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi). tan x + sec x = 1

OpenStudy (anonymous):

MY WORK: (sin x/ cos x) + (1/ cos x) = 1 (sin x + 1)/ cos x = 1 Then, I'm stuck. :(

Vocaloid (vocaloid):

subtract sec(x) from both sides first tan(x) = 1 - sec(x) then square both sides tan^2(x) = (1-secx)^2 = 1 - 2secx + sec^x then use the trigonometric identity tan^(x) = sec^2(x)-1 to change the left side sec^2(x) - 1 = 1 - 2secx + sec^x can you finish from here?

Vocaloid (vocaloid):

ah, a bit of a typo last expression should be sec^2(x) - 1 = 1 - 2secx + sec^2(x)

OpenStudy (anonymous):

So what I was doing is wrong?

Vocaloid (vocaloid):

well, not "wrong" but it would be difficult to get an answer by converting tan(x) and sec(x) into sin(x) and cos(x)

Vocaloid (vocaloid):

there might be a way to do it your way, but I can't think of an easy way to go from there

OpenStudy (anonymous):

Wait, so will it be equal to 0?

Vocaloid (vocaloid):

sec^2(x) - 1 = 1 - 2secx + sec^2(x) cancel out sec^2(x) -1 = 1 - 2sec(x) 2sec(x) = 2 sec(x) = 1 then solve for x

OpenStudy (anonymous):

Then it is 0. :D

Vocaloid (vocaloid):

yea